Find the equilibrium point for the demand and supply function D(p) =87-8p and S(p) =48+5p
Answer by robertb(5830)

The equilibrium point is where D(p) = S(p).
==> let 87 - 8p = 48 + 5p
==> 39 = 13p ==> p = 3
Therefore the equilibrium point is (3, 63). (Both supply and demand at 48 units at the price of 3.)